# Black hole accretion of scalar clouds with spontaneous symmetry breaking

### v1

@@ Abstract
Introduction. The past decade has witnessed spectacular progress in the field of black hole physics, driven primarily by the detection of gravitational waves from binary black hole mergers, which opened up new avenues for exploring the properties of black holes (Abbott:2016blz, ). Prime among these is the testing of the validity of the black hole uniqueness theorems of general relativity (see (Chrusciel:2012jk, ) for a review), particularly through the measurement of quasinormal modes during the ringdown phase (Price:2017cjr, ; Giesler:2019uxc, ). The Event Horizon Telescope has similarly revolutionized our understanding of black holes by capturing the images of M87* and Sgr A* (EventHorizonTelescope:2019dse, ; EventHorizonTelescope:2022wkp, ), revealing a striking feature: a luminous ring encircling a dark shadow. These distinctive signatures have been attributed to the intense light deflection occurring near unstable bound photon orbits, providing insights into black hole physics in the strong field regime, e.g. the structure of luminous accretion disks (Yuan:2014gma, ; Garcia:2013lxa, ; McKinney:2012vh, ; EventHorizonTelescope:2021srq, ; Blaes:2013toa, ).
@@
Scalarization in the EMS model occurs dynamically through a tachyonic (i.e. negative effective potential) instability of an initial scalar field perturbation on the background of a hairless black hole, driving the system into a new configuration with lower energy accompanied by a non-trivial scalar field distribution. In this Letter we ignore spin, so the initial state is a Reissner-Nordström (RN) black hole; notice that electric charge is needed to trigger the destabilization effect (Herdeiro:2018wub, ). The appearance of a nontrivial scalar profile, or condensate, may be naturally understood as a process of spontaneous symmetry breaking. In the original setup, the scalar field is real, so the symmetry that is broken by the scalarization effect is $\mathbb{Z}_{2}$, or $\Phi\mapsto-\Phi$. Here we focus instead on the case of a complex scalar as it provides a more interesting symmetry breaking pattern, namely that of a (global) $U(1)$ symmetry. The study of spontaneous scalarization in this scenario has been pioneered in (Latosh:2023cxm, ; Hyun:2024sfv, ).
In this Letter, we employ numerical relativity to investigate the fully nonlinear dynamics of the EMS model, specifically the spontaneous scalarization of a charged black hole and the subsequent evolution of the scalar cloud. After its initial formation following the tachyonic destabilization, we observe that the cloud continues to evolve, on a timescale of order the distance between the cloud and the event horizon, in a process of energy loss both to accretion into the black hole and ejection to spatial infinity. We find essential differences between the scenario where scalarization occurs starting in a symmetric phase, i.e. without further instabilities upon formation of a scalar condensate, versus the case where a scalar cloud forms from a symmetry-broken phase. Unlike with scalarization from the symmetric phase, the breaking of $U(1)$ symmetry allows for multiple accretion channels for the cloud, each leading to different amounts of energy loss and distinct final states. Notably, a dominant accretion channel emerges due to symmetry breaking.
Scalarization. We consider the EMS model of (Herdeiro:2018wub, ), extended here to a complex scalar field $\Phi=\phi_{1}+i\phi_{2}$, which is non-minimally coupled to the electromagnetic field $A_{\mu}$. The complete action is (we use geometrized units with $G=c=1$)
@@
where $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$. In this Letter, we adopt $f\left(\left|\Phi\right|\right)=e^{\alpha|\Phi|^{2}}$ for the non-minimal coupling function, with $\alpha$ a real constant, a choice which admits an electrovacuum solution given by a RN black hole and $\Phi=0$. Note that we do not include a minimal coupling of the scalar field to electromagnetism, i.e. $\Phi$ is uncharged under the electromagnetic $U(1)$, unrelated to the global $U(1)$ symmetry $\Phi\mapsto e^{i\varphi}\Phi$. The electrovacuum solution preserves the $U(1)$ symmetry, so we refer to it as the symmetric phase.
@@
More in detail, and considering now a general static and spherically symmetric ansatz for the metric, $ds^{2}=-e^{-2\delta(r)}N(r)dt^{2}+dr^{2}/N(r)+r^{2}d\Omega^{2}$, one finds the equation $\frac{d^{2}\Psi}{dx^{2}}+\left(\omega^{2}-V_{\rm eff}(x)\right)\Psi=0$ obeyed by the spherical, or monopole, perturbation $\Psi\equiv r\delta\Phi$, where $\omega$ is the frequency, $x$ is defined by $dx/dr\equiv e^{\delta(r)}/N(r)$ and $V_{\rm eff}$ is the effective potential.222 Explicitly, $V_{\rm eff}=\frac{N}{r^{2}e^{2\delta}}\left[1-N-2r^{2}\left|\Phi^{\prime}\right|^{2}-\frac{Q^{2}\left(1+\alpha-2\left|\alpha\Phi-r\Phi^{\prime}\right|^{2}\right)}{r^{2}e^{\alpha\left|\Phi\right|^{2}}}\right].$ In Fig. 1, the effective potential for the $U(1)$-symmetric phase is seen to display a negative well near the event horizon, and a numerical calculation shows that a tachyon mode with $\omega=0.2454iM$ is present. In agreement with expectations, this mode is responsible for triggering spontaneous scalarization (Herdeiro:2018wub, ; Guo:2024cts, ). Intriguingly, we find that there exist $U(1)$-broken phases (scalarized black holes with $\Phi\neq 0$), which also exhibit a negative potential well, a feature that is linked to the presence of a node in the scalar profile. This suggests that these symmetry-breaking phases are unstable to the formation of a scalar cloud localized around the potential well. Our aim in this Letter is to advance the investigation of the nonlinear dynamics of scalar clouds to encompass both $U(1)$-symmetric and $U(1)$-broken phases.
@@
Figure 2: Dynamical formation of a scalar cloud for the initial perturbation with $\theta_{0}=\pi/2$ in the $U(1)$-symmetric ($M=2.5$, leftmost column) and $U(1)$-broken ($\Phi_{h}=0.19$, right columns) phases.
@@
Figure 3: Dynamical evolution of the field norm $\left|\Phi\right|$ for initial perturbations with $\theta_{0}=0$ (upper row) and $\theta_{0}=\pi$ (lower) in the symmetry-broken phase ($\Phi_{h}=0.19$). The scalar cloud formed from the positive perturbation ($\theta_{0}=0$) can induce nonlinear effects at late times, causing the scalar field node to propagate towards spatial infinity. Conversely, the negative perturbation ($\theta_{0}=\pi$) leads to the node of the field moving into the event horizon at late times, resulting in a stable, nodeless scalar distribution outside the black hole.
Figure 4: Channels of scalar cloud accretion, represented by the value of the scalar field at the event horizon, $\Phi_{h}$. Upper row: In the $U(1)$-symmetric phase ($\Phi=0$ and $M=2.5$), the accretion dynamics and final states remain $U(1)$-symmetric with respect to the initial perturbation phase $\theta_{0}$. Lower: In the $U(1)$-broken phases, using $\Phi_{h}=0.1$ (left panel) and $\Phi_{h}=0.19$ (right), the accretion paths no longer respect the $U(1)$ symmetry. Cloud accretion tends to follow a dominant path, a phenomenon that becomes more pronounced for larger $\Phi_{h}$ values of the initial phase: in the right panel, the black and colored lines almost completely overlap.
@@
Figure 5: Upper panel: Energy of initial (dashed lines) and final (solid) states. Energy loss in the $U(1)$-symmetric phase is illustrated by the purple lines, and we consider two $U(1)$-breaking models (orange and blue). In the latter cases, the colored bands represent the range of final values of the energy. The bands are bounded by the cases with $\theta_{0}=0$ (squares) and $\theta_{0}=\pi$ (circles). The colored and black solid lines match those represented in Fig. 4. The zoomed-in insets allow one to discern the energy gap between initial and final states. Lower: The table provides explicit data for both the initial and end states in our numerics, in particular the energy loss. Here, $M_{0}$ represents the energy of the initial state, $M_{1}$ is the energy of the end state, and $\Delta M_{1}$ denotes the width of the energy band for the final states.
Energy loss. A noteworthy phenomenon during scalar cloud accretion is that the cloud reaches equilibrium after a portion of its energy is radiated to spatial infinity (see Figs. 2 and 3), potentially causing the final state to have less energy than the initial state. In Fig. 5, we calculate the Arnowitt-Deser-Misner mass (Misner:1973prb, ) of the initial and final states to quantify the energy loss during accretion in both the $U(1)$-symmetric and $U(1)$-breaking phases. In both cases, a discernible energy gap exists between the initial and final states, confirming that the formation of a stable scalar cloud involves the radiation of energy to spatial infinity.
@@
Conclusions. In the EMS model, the non-minimal coupling between the complex scalar and electromagnetic fields can induce a negative potential well near a RN black hole ($U(1)$-symmetric phase), triggering the onset of scalarization and leading to the accretion of a scalar cloud onto the event horizon. During dynamical evolution, the accretion of scalar clouds preserves the global $U(1)$ symmetry, resulting in $U(1)$-symmetric final states with energy that is degenerate with respect to the initial perturbation phase $\theta_{0}$. Remarkably, even in the $U(1)$-breaking phase, a negative potential well may exist in the presence of a non-trivial scalar field. Perturbations with different initial phases can induce the formation of scalar clouds within this well, which subsequently accrete through the event horizon, bringing energy to the black hole and radiating some energy to spatial infinity. Unlike the symmetry-preserving case, symmetry breaking allows for multiple accretion channels for scalar clouds with different initial phases. This phenomenon leads to varying amounts of energy radiated to spatial infinity, resulting in a lifting of degeneracy and the appearance of a band of possible final states. Moreover, symmetry breaking establishes a dominant accretion channel, acting as an attractor which draws in the majority of field trajectories.
This Letter highlights the significant impact of spontaneous symmetry breaking on the dynamics of scalar cloud accretion. Our studies reveal distinctive signatures of symmetry breaking in black holes, providing compelling motivation for their exploration in astrophysical observations. Future research could extend these findings to more realistic systems, such as rotating black holes, where the superradiance phenomenon is likely to produce richer scalar cloud structures and dynamics.

### v2

@@ Abstract
I Introduction
The past decade has witnessed spectacular progress in the field of black hole physics, driven primarily by the detection of gravitational waves from binary black hole mergers, which opened up new avenues for exploring the properties of black holes (Abbott:2016blz, ). Prime among these is the testing of the validity of the black hole uniqueness theorems of general relativity (see (Chrusciel:2012jk, ) for a review), particularly through the measurement of quasinormal modes during the ringdown phase (Price:2017cjr, ; Giesler:2019uxc, ). The Event Horizon Telescope has similarly revolutionized our understanding of black holes by capturing the images of M87* and Sgr A* (EventHorizonTelescope:2019dse, ; EventHorizonTelescope:2022wkp, ), revealing a striking feature: a luminous ring encircling a dark shadow. These distinctive signatures have been attributed to the intense light deflection occurring near unstable bound photon orbits, providing insights into black hole physics in the strong field regime, e.g. the structure of luminous accretion disks (Yuan:2014gma, ; Garcia:2013lxa, ; McKinney:2012vh, ; EventHorizonTelescope:2021srq, ; Blaes:2013toa, ).
@@ I Introduction
Scalarization in the EMS model occurs dynamically through a tachyonic (i.e. negative effective potential) instability of an initial scalar field perturbation on the background of a hairless black hole, driving the system into a new configuration with lower energy accompanied by a non-trivial scalar field distribution. In this paper we ignore spin, so the initial state is a Reissner-Nordström (RN) black hole; notice that electric charge is needed to trigger the destabilization effect (Herdeiro:2018wub, ). The appearance of a nontrivial scalar profile, or condensate, may be naturally understood as a process of spontaneous symmetry breaking. In the original setup, the scalar field is real, so the symmetry that is broken by the scalarization effect is $\mathbb{Z}_{2}$, or $\Phi\mapsto-\Phi$. Here we focus instead on the case of a complex scalar as it provides a more interesting symmetry breaking pattern, namely that of a (global) $U(1)$ symmetry. The study of spontaneous scalarization in this scenario has been pioneered in (Latosh:2023cxm, ; Hyun:2024sfv, ).
In this paper, we employ numerical relativity to investigate the fully nonlinear dynamics of the EMS model, specifically the spontaneous scalarization of a charged black hole and the subsequent evolution of the scalar cloud. After its initial formation following the tachyonic destabilization, we observe that the cloud continues to evolve, on a timescale of order the distance between the cloud and the event horizon, in a process of energy loss both to accretion into the black hole and ejection to spatial infinity. We find essential differences between the scenario where scalarization occurs starting in a symmetric phase, i.e. without further instabilities upon formation of a scalar condensate, versus the case where a scalar cloud forms from a symmetry-broken phase. Unlike with scalarization from the symmetric phase, the breaking of $U(1)$ symmetry allows for multiple accretion channels for the cloud, each leading to different amounts of energy loss and distinct final states. Notably, a dominant accretion channel emerges due to symmetry breaking.
II Scalarization
We consider the EMS model of (Herdeiro:2018wub, ), extended here to a complex scalar field $\Phi=\phi_{1}+i\phi_{2}$, which is non-minimally coupled to the electromagnetic field $A_{\mu}$. The complete action is (we use geometrized units with $G=c=1$)
@@ II Scalarization
[…]
@@
More in detail, and considering now a general static and spherically symmetric ansatz for the metric, $ds^{2}=-e^{-2\delta(r)}N(r)dt^{2}+dr^{2}/N(r)+r^{2}d\Omega^{2}$, one finds the equation $\frac{d^{2}\Psi}{dx^{2}}+\left(\omega^{2}-V_{\rm eff}(x)\right)\Psi=0$ obeyed by the spherical, or monopole, perturbation $\Psi\equiv r\delta\Phi$, where $\omega$ is the frequency, $x$ is defined by $dx/dr\equiv e^{\delta(r)}/N(r)$ and $V_{\rm eff}$ is the effective potential. [^3] Explicitly, $V_{\rm eff}=\frac{N}{r^{2}e^{2\delta}}\left[1-N-2r^{2}\left|\Phi^{\prime}\right|^{2}-\frac{Q^{2}\left(1+\alpha-2\left|\alpha\Phi-r\Phi^{\prime}\right|^{2}\right)}{r^{2}e^{\alpha\left|\Phi\right|^{2}}}\right].$ In Fig. 1, the effective potential for the $U(1)$-symmetric phase is seen to display a negative well near the event horizon, and a numerical calculation shows that a tachyon mode with $\omega=0.2454iM$ is present. In agreement with expectations, this mode is responsible for triggering spontaneous scalarization (Herdeiro:2018wub, ; Guo:2024cts, ). Intriguingly, we find that there exist $U(1)$-broken phases (scalarized black holes with $\Phi\neq 0$), which also exhibit a negative potential well, a feature that is linked to the presence of a node in the scalar profile. This suggests that these symmetry-breaking phases are unstable to the formation of a scalar cloud localized around the potential well. Our aim in this paper is to advance the investigation of the nonlinear dynamics of scalar clouds to encompass both $U(1)$-symmetric and $U(1)$-broken phases.
III Dynamical formation of scalar clouds
@@
Figure 2: Dynamical formation of a scalar cloud for the initial perturbation with $\theta_{0}=\pi/2$ in the $U(1)$-symmetric ($M=2.5$, leftmost column) and $U(1)$-broken ($\Phi_{h}=0.19$ and $M=0.7744$, right columns) phases. A scalar cloud develops within the negative potential well due to scalarization and is subsequently accreted onto the black hole, terminating in the formation of a stable cloud structure around it.
@@
Figure 3: Dynamical evolution of the field norm $\left|\Phi\right|$ for initial perturbations with $\theta_{0}=0$ (upper row) and $\theta_{0}=\pi$ (lower) in the symmetry-broken phase ($\Phi_{h}=0.19$ and $M=0.7744$). The scalar cloud formed from the positive perturbation ($\theta_{0}=0$) can induce nonlinear effects at late times, causing the scalar field node to propagate towards spatial infinity. Conversely, the negative perturbation ($\theta_{0}=\pi$) leads to the node of the field moving into the event horizon at late times, resulting in a stable, nodeless scalar distribution outside the black hole.
Figure 4: Channels of scalar cloud accretion, represented by the value of the scalar field at the event horizon, $\Phi_{h}$. Upper row: In the $U(1)$-symmetric phase ($\Phi=0$ and $M=2.5$), the accretion dynamics and final states remain $U(1)$-symmetric with respect to the initial perturbation phase $\theta_{0}$. Lower: In the $U(1)$-broken phases, using here $\Phi_{h}=0.1$, $M=1.8744$ (left panel) and $\Phi_{h}=0.19$, $M=0.7744$ (right panel), the accretion paths no longer respect the $U(1)$ symmetry. Cloud accretion tends to follow a dominant path, a phenomenon that becomes more pronounced for larger $\Phi_{h}$ values of the initial field: in the right panel, the black and colored lines almost completely overlap (the visible purple and orange lines cover all other colored as well as black lines). As a result, the final states are not $U(1)$-symmetric, as indicated by the varying values of $\left|\Phi_{h}\right|$ shown in the lower right corners.
@@
The existence of a dominant path in field space may be understood from the fact that one expects most of the energy of the field to go into the gapless degrees of freedom, in this case the angular component of $\Phi$, with a comparatively smaller fraction in the radial field. [^6] We would like to thank an anonymous referee for raising this interesting question. This picture is consistent with the observed trajectory, which appears to be roughly circular, at least qualitatively. This intuition is corroborated by an explicit comparison of the energy densities stored in the radial and angular components of the field; see Appendix B for details.
Figure 5: Upper panel: Energy of initial (dashed lines) and final (solid) states. Energy loss in the $U(1)$-symmetric phase is illustrated by the purple lines, and we consider two $U(1)$-breaking models (orange and blue). In the latter cases, the colored bands represent the range of final values of the energy. The bands are bounded by the cases with $\theta_{0}=0$ (squares) and $\theta_{0}=\pi$ (circles). The zoomed-in insets allow one to discern the energy gap between initial and final states. Lower: The table provides explicit data for both the initial and end states in our numerics, in particular the energy loss. Here, $M_{0}$ represents the energy of the initial state, $M_{1}$ is the energy of the end state, and $\Delta M_{1}$ denotes the width of the energy band for the final states.
IV Energy loss
A noteworthy phenomenon during scalar cloud accretion is that the cloud reaches equilibrium after a portion of its energy is radiated to spatial infinity (see Figs. 2 and 3), potentially causing the final state to have less energy than the initial state. In Fig. 5, we calculate the Arnowitt-Deser-Misner mass (Misner:1973prb, ) of the initial and final states to quantify the energy loss during accretion in both the $U(1)$-symmetric and $U(1)$-breaking phases. In both cases, a discernible energy gap exists between the initial and final states, confirming that the formation of a stable scalar cloud involves the radiation of energy to spatial infinity.
@@ IV Energy loss
V Conclusions
In the EMS model, the non-minimal coupling between the complex scalar and electromagnetic fields can induce a negative potential well near a RN black hole ($U(1)$-symmetric phase), triggering the onset of scalarization and leading to the accretion of a scalar cloud onto the event horizon. During dynamical evolution, the accretion of scalar clouds preserves the global $U(1)$ symmetry, resulting in $U(1)$-symmetric final states with energy that is degenerate with respect to the initial perturbation phase $\theta_{0}$. Remarkably, even in the $U(1)$-breaking phase, a negative potential well may exist in the presence of a non-trivial scalar field. Perturbations with different initial phases can induce the formation of scalar clouds within this well, which subsequently accrete through the event horizon, bringing energy to the black hole and radiating some energy to spatial infinity. Unlike the symmetry-preserving case, symmetry breaking allows for multiple accretion channels for scalar clouds with different initial phases. This phenomenon leads to varying amounts of energy radiated to spatial infinity, resulting in a lifting of degeneracy and the appearance of a band of possible final states. Moreover, symmetry breaking establishes a dominant accretion channel, acting as an attractor which draws in the majority of field trajectories.
This paper highlights the significant impact of spontaneous symmetry breaking on the dynamics of scalar cloud accretion. Our studies reveal distinctive signatures of symmetry breaking in black holes, providing compelling motivation for their exploration in astrophysical observations. The assumption of spherical symmetry in our setup is admittedly simplistic, and implies that neither electromagnetic nor gravitational radiation will be emitted during the evolution. It would be therefore interesting to relax this assumption in order to investigate these radiation channels and how the properties of the system may be assessed through their observation. Future research could also extend these findings to more realistic systems, such as rotating black holes, where the superradiance phenomenon is likely to produce richer scalar cloud structures and dynamics.
@@ V Conclusions
Appendix A Numerical scheme for black hole evolution
In this paper, we simulate the black hole evolution using the $3+1$ decomposition of the metric, expressed as
$ds^{2}=-N_{0}^{2}dt^{2}+\gamma_{ij}\left(dx^{i}+N^{i}dt\right)\left(dx^{j}+N^{j}dt\right)\,.$ (2)
For the gravity sector, we apply the BSSN formulation, using the 1+log slicing and Gamma-driver shift conditions (Bona:2003fj, ; Brown:2009dd, ; Etienne:2014tia, ; Ruchlin:2017com, ; Etienne:2024ncu, ; Baumgarte:2012xy, ).
To model the scalar field, we employ dynamical variables $\Phi$ and $\Pi$, where $\Pi=n^{\mu}\nabla_{\mu}\Phi$, where $n_{\mu}$ is given by $n_{\mu}=\left(-N_{0},0,0,0\right)$, as the 4-vector orthogonal to the spatial hypersurface. For the electromagnetic field, we decompose $F_{\mu\nu}$ into the 3-dimensional electric and magnetic components, defined as $E_{i}=\gamma_{i}^{\mu}n^{\nu}F_{\mu\nu}$ and $B_{i}=\gamma_{i}^{\mu}n^{\nu}*F_{\mu\nu}$, respectively, where * denotes the Hodge dual, and $\gamma_{i}^{\mu}$ projects 4-dimensional vectors onto the spatial hypersurface. Since we focus on spherically symmetric evolution, the magnetic field $B_{i}$ vanishes during the simulation. The evolution equations for the matter fields read
$\partial_{\perp}\Phi$ $=N_{0}\Pi\,,$
$\partial_{\perp}\Pi$ $=D_{i}\left(N_{0}\chi^{i}\right)+N_{0}\Pi K-\frac{1}{2}\alpha N_{0}\Phi e^{\alpha|\Phi|^{2}}F^{2}\,,$
$\partial_{\perp}E^{i}$ $=\alpha KE^{i}-2\alpha N_{0}E^{i}\left(\phi_{1}\Pi_{1}+\phi_{2}\Pi_{2}\right)\,,$ (3)
where the time derivative is defined by $\partial_{\perp}\equiv\partial_{t}-\mathcal{L}_{N}$, and $\mathcal{L}_{N}$ is the Lie derivative along $N_{i}$; $K$ is the trace of the extrinsic curvature, $F^{2}=-2E^{i}E_{i}$, and $\phi_{1,2}$ and $\Pi_{1,2}$ denote respectively the real and imaginary parts of $\phi$ and $\Pi$.
Figure 6: Convergence test of the numerical scheme, illustrated by the event-horizon value $|\Phi_{h}|$ of the scalar field as function of the number $N_{R}$ of grid points. The blue dots represent the final state of spontaneous scalarization from a RN black hole with $Q/M=0.5$ and $\alpha=160$. As the grid number increases, $|\Phi_{h}|$ converges, indicating convergence of the simulation.
During the evolution, the gravity-matter interaction is described by the stress-energy tensor $T_{\mu\nu}$, given by
$T_{\mu\nu}$ $=\frac{1}{8\pi}\left(2\partial_{(\mu}\Phi^{*}\partial_{\nu)}\Phi-g_{\mu\nu}\left|\partial\Phi\right|^{2}\right)$
$\quad+\frac{1}{8\pi}e^{\alpha|\Phi|^{2}}\left(2F_{\mu\rho}F_{\nu}{}^{\rho}-\frac{1}{2}g_{\mu\nu}F^{2}\right)\,.$ (4)
To incorporate the matter contribution into the BSSN formalism, we project the stress-energy tensor onto $3+1$ variables,
$\rho$ $=n^{\mu}n^{\nu}T_{\mu\nu}=\frac{1}{8\pi}\gamma^{ij}\left(\partial_{i}\phi_{1}\partial_{j}\phi_{1}+\partial_{i}\phi_{2}\partial_{j}\phi_{2}\right)$
$\quad+\frac{1}{8\pi}\left(\Pi_{1}\Pi_{1}+\Pi_{2}\Pi_{2}e^{\alpha|\Phi|^{2}}E^{2}\right)\,,$
$J_{i}$ $=-\gamma_{i}^{\mu}n^{\nu}T_{\mu\nu}=-\frac{1}{4\pi}\left(\partial_{i}\phi_{1}\Pi_{1}+\partial_{i}\phi_{2}\Pi_{2}\right)\,,$
$S_{ij}$ $=\gamma_{i}^{\mu}\gamma_{j}^{\nu}T_{\mu\nu}$
$=\frac{1}{4\pi}\left(\partial_{i}\phi_{1}\partial_{j}\phi_{1}+\partial_{i}\phi_{2}\partial_{j}\phi_{2}-e^{\alpha|\Phi|^{2}}E_{i}E_{j}\right)$
$\quad-\frac{1}{8\pi}\gamma_{ij}\gamma^{kl}\left(\partial_{k}\phi_{1}\partial_{l}\phi_{1}+\partial_{k}\phi_{2}\partial_{l}\phi_{2}\right)$
$\quad+\frac{1}{8\pi}\gamma_{ij}\left(\Pi_{1}\Pi_{1}+\Pi_{2}\Pi_{2}+e^{\alpha|\Phi|^{2}}E^{2}\right)\,.$ (5)
In our numerical simulations, we adopt a spherical-like coordinate system, introducing a dimensionless radial coordinate $R$ to replace the standard spherical radial coordinate $r$. The relationship between $r$ and $R$ is defined as
$r=r_{\rm max}\left(R_{0}+\frac{e^{R/a}-e^{-R/a}}{e^{1/a}-e^{-1/a}}\right)\,,$ (6)
where $r_{\rm max}$ denotes the outer boundary, and $R_{0}$ and $a$ are scaling constants that map between $r$ and $R$. Since we focus on a spherically symmetric spacetime, the radial direction $R$ is discretized uniformly with $N_{R}$ grid points, while $N_{\theta}=N_{\varphi}=1$ for the angular directions $\left(\theta,\varphi\right)$.
In Fig. 6, we present a convergence test for the simulation of spontaneous scalarization of a RN black hole, evaluated across different grid resolutions $N_{R}$. In the simulation setup, the radial coordinate parameters are specified as $r_{\rm max}=30000$, $r_{0}=0.00012$ and $a=0.07$ (recall that we use units such that $Q=1$). The integration time step $\Delta t$ is chosen to match the smallest spatial $\Delta r_{\rm min}$ (i.e., $\Delta t=\Delta r_{\rm min}$) to avoid numerical instabilities. Our numerical results indicate that the simulation converges with increasing grid number $N_{R}$.
Appendix B Radial-angular scalar field decomposition
In this Appendix we expand on the relationship between the dominant path observed in the scalar cloud accretion channels and the question of energy distribution among the field components. At a qualitative level, and at first approximation, the attractor trajectory in Fig. 4 (black curves in bottom left panel) appears to correspond to circular motion, suggesting that the energy density of the scalar field is predominantly in the angular field. This agrees with the expectation that most of the energy should go to the gapless degrees of freedom in the system, with a comparatively smaller amount going to the radial part of the scalar field.
We can confirm this intuition through an explicit calculation of the energy density of each field component, with the radial and angular fields defined via
$\Phi=\sigma(x)e^{i\theta(x)}\,.$ (7)
We define the stress-energy tensors associated to $\sigma$ and $\theta$ by
$T^{(\sigma)}_{\mu\nu}$ $\equiv T_{\mu\nu}\Big{|}_{\theta=\theta_{0},A=0}$
$=\frac{1}{8\pi}(2\partial_{\mu}\sigma\partial_{\nu}\sigma-g_{\mu\nu}\partial^{\rho}\sigma\partial_{\rho}\sigma)\,,$
$T^{(\theta)}_{\mu\nu}$ $\equiv T_{\mu\nu}\Big{|}_{\sigma=\sigma_{0},A=0}$
$=\frac{1}{8\pi}(2\sigma_{0}^{2}\theta^{2}\partial_{\mu}\theta\partial_{\nu}\theta-g_{\mu\nu}\theta^{2}\sigma_{0}^{2}\partial^{\rho}\theta\partial_{\rho}\theta)\,,$ (8)
and the energy density of each component is given by $\rho_{a}=n^{\mu}n^{\nu}T^{(a)}_{\mu\nu}$ (cf. (5)).
The time dependence of $\rho_{\sigma}$ and $\rho_{\theta}$ is shown in Fig. 7 for the same simulations considered in the main text (cf. in particular Fig. 5). The results make it clear that the peak energy density of the angular field dominates over the peak value for the radial field, at least by an order of magnitude. Also noteworthy is the fact that the peak in $\rho_{\theta}$ has a relatively short duration, compared to the typical timescale for the accretion process. This is further illustrated in Fig. 8, which shows that most of the field trajectory is traversed in the comparatively short time interval during which the angular field dominates the energy budget.
Figure 7: Time evolution of the energy densities $\rho_{\sigma}$ and $\rho_{\theta}$, evaluated at the event horizon, and using the initial values $\sigma_{0}=0.19$ and $\sigma_{0}=0.10$. One observes that both the radial and angular fields experience a peak in the energy density corresponding to the process of accretion onto the black hole.
Figure 8: Trajectory of the horizon value of the scalar field, $\Phi_{h}$, in the $\phi_{1}-\phi_{2}$ plane, using the initial values $\sigma_{0}=0.19$ and $\sigma_{0}=0.10$. Dots are separated by equal time intervals, and their color indicates the ratio of $\rho_{\theta}$ to its maximum value.